# Diversity, equity, and inclusion for problems in additive number theory

### 16 Mar 2026 v1

@@ Abstract.
This is a survey the diversity of problems in additive number theory. Equity requires the consideration of less currently popular problems, and suggests their inclusion in the additive canon. Of particular interest are problems about the sizes of sumsets of finite sets of integers and problems about the arithmetical structure of intersections of sumsets.
@@ 1. Canonical additive problems
These are the canonical additive problems. (Waring and Goldbach were discussed in my book Additive Number Theory: The Classical Bases [12].) I shall ignore these topics and, instead, focus on other less popular and more “diverse” questions that have not yet been or have not always been included in the additive canon. Fairness, or equity, suggests that less well-known and not obviously “important” topics must always be considered candidates for inclusion, because only time and not today’s “experts” will decide what is really important.
@@ 2. Sum-product: A problem no one cared about
$f_{2}(n)>\frac{n^{2}}{(\log n)^{k}}$
is certainly false.
In 1983, Erdős and Szemerédi [5] proved that there exists an $\varepsilon>0$ such that
@@
$\varepsilon=1/31\operatorname{\qquad\text{and}\qquad}f_{2}(n)>cn^{32/31}$
@@ 4. A noncanonical problem: Sizes of sumsets
The lower bound comes from $k$-term arithmetic progressions: If
$a_{i},a_{j}\in\{0,1,2,\ldots,k-1\}$
then
$a_{i}+a_{j}\in\{0,1,2,\ldots,2k-2\}.$
The upper bound is combinatorial: There are $\binom{k}{2}$ distinct sums $a_{i}+a_{j}$ from a 2-element set and $k$ sums $a_{i}+a_{i}$, and
@@
Let
$2A=\{a+a^{\prime}:a,a^{\prime}\in A\}$
and
@@
$\mathcal{R}(2,k)=\left\{\left|2A\right|:A\subseteq\mathbf{Z}\text{ and }|A|=k\right\}$
@@
For all positive integers $k$, let
@@ 5. Range of sumset sizes
The $h$-fold sumset of a set $A$ of integes is the set of all sums of $h$ not necessarily distinct elements of $A$:
@@ 6. Missing sumset sizes for 3-fold sums of sets with 3 elements
$\mathcal{R}(3,k)=\{3k-2\}\cup\left[3k,\binom{3k+2}{3}\right$].
The original missing sumset size result was the following:
@@
For all $k$ there exists $h_{k}$ such that, for all $h>h_{k}$,
@@ 8. Sumset races
By theorems of Nathanson (for integers) and Khovanskii, (for groups), one cannot prescribe infinite sequences of $n$-tuples in sumset races.
@@
Theorem 11 (Kravitz).
Let $\tau_{1},\tau_{2},\ldots,\tau_{H}$ be $n$-tuples of positive integers. There exist finite sets $A_{1},\ldots,A_{n}$ such that the sumset size $n$-tuple
$(|hA_{1}|,|hA_{2}|,\ldots,|hA_{n}|)$
has the same relative order as the $n$-tuple $\tau_{h}$ for all $h=1,\ldots,H$.
There is a continuous analogue of this result.
Theorem 12 (Nathanson).
@@ 10. Sumset intersection limits
If $A$ is a set of nonnegative integers and if $(A_{q})_{q=1}^{\infty}$ is a strictly decreasing sequence of sets of nonnegative integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$, then $\mathcal{H}(A_{q})=\mathbf{N}$.
@@
for all asymptotically decreasing sequences in an additive abelian semigroup. It is not even true in the integes. Diego Marques and Nathanson proved the following:
@@ 11. The moral of the story
It suggests one of Aesop’s fables. Aesop lived in ancient Greece between 620 and 564 BCE. Herodotus, in the fifth century BCE, mentioned him in his Histories in Book II, Chapter 134, where he referred to him as a fable writer and says he was a slave…. Aesop’s fables were transmitted orally and not collected and written down until three centuries after Aesop’s death. The fables were part of an oral tradition and were not collected until about three centuries after Aesop’s death. Stories have been added continuously and are still being added, so it is not clear which of Aesop’s fables were recited by Aesop.
@@
This is a good way to think about problems in additive number theory.

### 19 Mar 2026 v2

@@ Abstract.
This is a survey of the diversity of problems in additive number theory. Equity requires the consideration of less currently popular problems, and suggests their inclusion in the additive canon. Of particular interest are problems about the sizes of sumsets of finite sets of integers and problems about the arithmetical structure of intersections of sumsets.
@@ 1. Canonical additive problems
These are the canonical additive problems. (Waring and Goldbach were discussed in my book Additive Number Theory: The Classical Bases [12].) I shall ignore these topics and, instead, focus on other less popular and more “diverse” questions that have not yet been or have not always been included in the additive canon. Fairness, or equity, suggests that less well-known and not obviously “important” topics must always be considered candidates for inclusion, because only time and not today’s “experts” decide what is really important.
@@ 2. Sum-product: A problem no one cared about
Several years later, in 1983, Erdős and Szemerédi [5] proved that there exists an $\varepsilon>0$ such that
@@
$\varepsilon=\frac{1}{31}\operatorname{\qquad\text{and}\qquad}f_{2}(n)>cn^{32/31}$
@@ 4. A noncanonical problem: Sizes of sumsets
The lower bound comes from $k$-term arithmetic progressions. Every integer in the interval of integers $\{0,1,2,\ldots,2k-2\}$ is a sum of two integers in the arithmetic progression $A=\{0,1,2,\ldots,k-1\}$. Moreover, the set of 2-fold sums of a set of $k$ integers has size 2k12k-1 if and only if the set is an arithmetic progression.
The upper bound is combinatorial: If $A=\{a_{1},\ldots,a_{k}\}$ is a set of size $k$ whose $\binom{k}{2}$ sums $a_{i}+a_{j}$ with $a_{i}\neq a_{j}$ and $k$ sums $a_{i}+a_{i}$ are distinct, then the number of 2-fold sums is
@@
For every set $A$ of integers, let
$2A=\{a+a^{\prime}:a,a^{\prime}\in A\}.$
@@
$\mathcal{R}(2,k)=\left\{\left|2A\right|:A\subseteq\mathbf{Z}\text{ and }|A|=k\right\}.$
@@
For every set $A$ of integers, let
For all positive integers $k$, we define
@@ 5. Range of sumset sizes
The $h$-fold sumset of a set $A$ of integers is the set of all sums of $h$ not necessarily distinct elements of $A$:
@@ 6. Missing sumset sizes for 3-fold sums of sets with 3 elements
$\mathcal{R}(3,k)=\{3k-2\}\cup\left[3k,\binom{k+2}{3}\right$].
Problem 6.
For all $k\in\mathbf{N}$,
$\mathcal{R}(4,k)\subseteq\left[4k-3,\binom{k+3}{4}\right$].
Compute $\mathcal{R}(4,k)$ for all $k\in\mathbf{N}$.
The firsrt missing sumset size result was the following:
@@
For all $h$ there exists $k_{h}$ such that, for all $k>k_{h}$,
@@ 8. Sumset races
By theorems of Nathanson (for integers) and Khovanskii (for groups), one cannot prescribe infinite sequences of $n$-tuples in sumset races.
@@
There is a continuous analogue of Theorem 10.
Theorem 11 (Nathanson [20]).
@@ 10. Sumset intersection limits
If $A$ is a set of nonnegative integers and if $(A_{q})_{q=1}^{\infty}$ is a strictly decreasing sequence of sets of nonnegative integers such that $A=\bigcap_{q=1}^{\infty}A_{q}$, then
$hA=\bigcap_{q=1}^{\infty}hA_{q}$
for all $h=1,2,3,\ldots$, that is,
$\mathcal{H}(A_{q})=\mathbf{N}.$
@@
for all asymptotically decreasing sequences in an additive abelian semigroup. It is not even true in the integers. Diego Marques and Nathanson proved the following:
@@
Problem 16.
Compute the set
$\mathcal{H}^{*}(0)=\left\{\mathcal{H}(A_{q}):\bigcap_{q=1}^{\infty}A_{q}=\{0\}\right\}.$
@@ 11. The moral of the story
It suggests one of Aesop’s fables. Aesop lived in ancient Greece between 620 and 564 BCE. Herodotus, in the fifth century BCE, mentioned him in his Histories in Book II, Chapter 134, where he referred to him as a fable writer and says he was a slave…. Aesop’s fables were transmitted orally and not collected and written down until three centuries after Aesop’s death. Stories have been added continuously and are still being added, so it is not clear which of Aesop’s fables were recited by Aesop.
@@
This is a good way to think about the diversity of problems in additive number theory.